### v1

@@ 2.1.4 Bias toward different patient groups
The problem of bias in machine learning model performance in medicine is a important concern as it impacts the reliability and fairness of medical decisions, potentially leading to suboptimal care and health disparities among different population subgroups [22, 23, 24]. The development and deployment of AI in medicine faces several bias-related challenges: (1) Label Bias. The labels in training data might reflect existing biases in clinical practice, mirroring the unequal attention and care provided to different demographic groups, thereby influencing AI outcomes. (2) Algorithmic Bias. The choice of algorithms, features, and model parameters can introduce and amplify biases. This affects model predictions and outcomes, potentially leading to inequitable healthcare delivery. (3) Evaluation Bias. Evaluation metrics and benchmarks might be biased towards certain groups or conditions. This can mask the true performance of AI models across diverse populations. (4) Socioeconomic and Cultural Bias. Differences in access to healthcare resources, healthcare-seeking behaviors, and disease prevalence across socioeconomic and cultural groups can lead to bias in construction of AI models. (5) Historical Bias. Historical inequalities and prejudices in medical research and clinical practice can be perpetuated and amplified through biased machine learning models, necessitating a conscientious approach to model training and validation.
2.1.5 Difficulty in adapting to a new patient population
@@
Fine-tuning machine learning models to local data in medicine is essential to adapting models to the specific characteristics of new patient populations and healthcare settings. However, it poses several challenges that impact the effectiveness and generalizability of models. [25, 26] Apart from the Regulatory and Ethical Issues discussed before and the aforementioned data related challenges including Data Scarcity, Data Quality and Variability and Heterogeneity, Computational Resources may also lead to safety issues, given that the limitation of computational resources in local environments can be a significant barrier to the fine-tuning of complex models. Moreover, Domain Expertise can be a problem as well, as the absence of local expertise in fields like machine learning and medicine can pose challenges to the adaptation and validation of models.
2.2 Alignment problems of artificial intelligence in medicine
@@
Besides performing different tasks reliably and robustly, it is important to ensure the AI models’ behaviors reflect human users’ goals and values. In this section, we discuss the challenges of alignment problems for AI in medicine, with a focus on technical and normative alignment as well as AI-clinician Interaction.
2.2.1 Alignment challenges when using AI in medical practice
@@
The alignment problem in AI refers to the challenge of ensuring that artificial intelligence systems reliably behave in ways that are in accordance with human values, intentions, and expectations. This problem arises from the inherent difficulty in specifying complex human values and objectives precisely and completely in a manner that AI systems can understand and implement. The goal of AI value alignment is to ensure that AI is properly aligned with human values [27].
@@
The challenge of alignment has two parts. The first is technical, focusing on the formal incorporation of values or principles into AI to ensure their reliable and appropriate actions. Real-world instances of AI misalignment have been observed, such as chatbots promoting abusive content when interacting freely online [28]. Moreover, this challenge escalates with more advanced AIs, raising issues like ‘reward-hacking’—where agents find unanticipated methods to fulfill objectives, diverging from intended outcomes—and the difficulty of evaluating AIs whose cognitive abilities might significantly surpass human capabilities [29].
@@
The second aspect of the value alignment problem is normative, raising the question of which values or principles should be encoded into AIs. This can be viewed through two lenses: minimalist and maximalist conceptions of value alignment. The minimalist approach seeks to bind AI to a viable framework of human values to prevent unsafe outcomes, while the maximalist perspective aims for alignment with the most accurate or optimal human values on a broader societal or global scale. Although the minimalist perspective rightly notes that optimizing for nearly any single metric could yield adverse outcomes for humans, advancing beyond this approach might be necessary for achieving full alignment with AI. This is because AI systems could be safe and reliable but still fall short of optimal or truly desirable outcomes [27]. Here we conclude the alignment challenges as in Table 3.
@@
Scalability and Complexity With increasing complexity and capability, ensuring that AI models remain aligned with human values and controlling their behavior becomes more challenging.
@@
2.2.2 AI-clinician interaction and alignment
@@
AI-clinician interaction and alignment is a fundamental aspect of implementing AI solutions in medical practice. A symbiotic collaboration between AI and clinicians, achieved through consensus finding or disagreements cross-validation, would significantly improve the quality of patient care and reduce clinical operation costs [30]. Several challenges need to be addressed to ensure effective and seamless integration of AI into clinical workflows. Complex and non-intuitive interfaces can impede effective interaction between clinicians and AI tools, adversely affecting adoption and user experience. Skepticism about the reliability of AI predictions may arise among clinicians, particularly when models lack transparency, interpretability, or validation in real-world settings. Incorporating AI tools into existing clinical workflows can be challenging, potentially causing disruptions and diminishing efficiency. Issues concerning patient privacy, data security, and accountability in cases of erroneous AI predictions can affect clinicians’ willingness to employ these tools. The performance of AI tools can be compromised by using inaccurate, incomplete, or outdated data, reducing their utility and reliability in clinical settings. A gap in AI understanding among clinicians and insufficient medical knowledge among AI developers can lead to less-than-optimal AI tool development and implementation. Inadequate training and education regarding AI tools can hinder clinicians’ ability to utilize and interpret outputs from them effectively. The integration of AI tools into clinical practice might alter the dynamics of patient-clinician relationships, potentially impacting trust, communication, and the overall quality of care. Advanced AI models, such as deep neural networks, often lack transparency, making it challenging to understand and anticipate their behaviors and underlying decision-making processes.
3 Risks of using AI in medicine
Recent years have seen a significant increase in the adoption of AI technologies in healthcare, affecting clinical practice, public health, and many other areas. While AI has brought promising improvements in efficiency and quality, it also presents potential risks in clinical practice, healthcare operations, and social impacts.
3.1 Safety issues in clinical practice
@@
In clinical practice, AI safety concerns exist in various forms, each carrying distinct risks. Although AI is usually designed and functions as assistive tools to support clinical decisions (e.g., diagnostic-AI, clinical monitoring, etc.[31, 32, 33]), unrobust, unsafe, and unexplainable black-box AI models could potentially result in disease misdiagnosis and subsequent mismanagement [34]. It’s been known that AI models with high predictive performance can be poorly calibrated and calibration-reporting guidelines, despite well adopted in the medical community [35], are rarely followed in the engineering research studies. Uncertainty measures of predicted outcomes, parameters, and statistics (e.g., confidence intervals) are also commonly under-reported in incomplete and/or inconsistent ways [36]. Poor generalization in AI systems can lead to misdiagnoses and inappropriate treatments, impacting patient health resource allocation and eroding trust in AI technologies. Biased models exacerbate these issues, particularly affecting underrepresented groups, leading to misdiagnoses, inappropriate treatments, and widening health disparities. Lastly, the challenge of fine-tuning of AI models using local data is essential to ensure equitable healthcare outcomes, underscoring the need for models sensitive to diverse patient populations and healthcare settings. Each of these risks highlights the critical need for rigorous testing, validation, and ethical considerations in the deployment of AI in medicine [37, 38, 39, 40].
3.2 Safety issues in healthcare system operation
@@
The integration of AI in healthcare systems, such hospital operations, community clinic operations, pharmacy operation and health insurance, brings various safety issues that could lead to risks and dangers. Data security and privacy is one of the largest issues when adopting AI in healthcare operations, as AI systems require access to vast amounts of sensitive patient data. There is a risk of data breaches, unauthorized access, or misuse of this data, which can lead to privacy violations and compromise patient confidentiality [41, 42]. Another issue is reliability, which refers to the possibility that AI systems might make errors in diagnosis or treatment recommendations due to limitations in their algorithms or training data. These errors can have serious consequences for patient health [43]. Ensuring that AI systems comply with existing healthcare regulations and standards is crucial. Non-compliance can lead to legal and ethical issues. Additionally, interoperability and system integration pose challenges in healthcare as integrating AI into existing systems can be complicated due to inter-operator discrepancies and differences in data standards, which can lead to inefficiencies and errors in patient care [12]. Resource allocation should be taken into consideration, as AI-driven decisions might influence the distribution of resources in healthcare, potentially prioritizing certain groups or treatments over others, which could exacerbate inequalities [44]. Apart from these technical problems, dependence and skill degradation among healthcare workers can occur due to over-reliance on AI, leading to a decrease in their decision-making skills [45].
3.3 Social risks of applying AI in medicine
@@
AI safety issues in medicine when presented on a societal scale, can lead to a range of broader risks and challenges. Here we outline general issues that may arise from policy level down to population level in table 2. These include public trust and perception, the ethical and societal norms, and the bias, fairness issues and more.
@@
Table 2: Social risks of applying AI in medicine.
@@
Widening Health Disparities If AI technologies are primarily accessible to well-resourced healthcare systems or populations, it could exacerbate existing health disparities. This has been observed in other technological advances and is termed the Matthew principle. Those without access to AI-enabled healthcare might receive lower quality care, widening the health outcome gap between different socio-economic groups [46].
@@
Policy and Regulatory Challenges As AI continues to advance, it may outpace current regulations and ethical guidelines, leading to legal and ethical dilemmas. Ensuring that policies keep up with technological advancements is crucial but challenging, and failure to do so could lead to societal risks [47].
@@
Misinformation and Manipulation Risks There’s a risk of AI being used to spread misinformation about health issues or being manipulated to provide misleading information, which can have widespread negative effects on public health and safety [48].
@@
Global Inequities The uneven distribution and adoption of AI in healthcare across different countries can exacerbate global health inequities. Countries with limited access to AI technology may fall behind in healthcare advancements [49].
@@
Dependency and System Vulnerabilities Over-reliance on AI systems can make healthcare infrastructure vulnerable to technical failures, cyberattacks, or other disruptions, potentially leading to widespread healthcare service interruptions [50].
4 AI safety issues related to large language models in medicine
Large language models (LLMs) like GPT-4-omni and Llama 3 have a wide range of applications in medicine. LLMs have been adapted for clinical documentation and reporting, medical information retrieval, disease prediction and diagnosis, and drug discovery and development and many other tasks [51, 52, 53, 54]. In recent years, a combination of a shortage of well-trained physicians and the increased complexity in the medical field constitutes a significant challenge for the timely and accurate delivery of healthcare. LLMs have been tested to diagnose complex clinical cases and show better diagnosis than human annotators [55]. Therefore, LLMs, together with machine learning models, could possibly be the key solutions to improve efficiency in healthcare. However, as a new interface between humans, data and computers, LLMs have many unique safety issues that are rarely found in other types of machine learning models. Over the past year leading up to the publication of this manuscript, we have witnessed a surge in research dedicated to enhancing the safety of LLMs from various angles. In this section, we aim to discuss different problems of LLMs that may lead to potential medical risks when applied in healthcare, as summarized in Figure 3.
It is pointed out that when using generative AI such as GPT-4 for medical purpose, two major safety problems are [56]: “Do not trust it. This type of AI makes things up when it does not know an answer. Never ever act on what it tells you without checking” and “You should know, though, that if you paste medical information into an online AI program, it loses our privacy protections”, The first point is a problem referred to as “hallucination” in the AI community. The second point is the privacy issue of sensitive medical data.
Figure 3: AI safety issues in medicine related to large language models.
Various limitations of inherent functionality of current LLMs raise safety concerns in medical applications. From issues of data accuracy and privacy to the models’ constraints in understanding complex medical data, we critically examine the facets that must be navigated carefully to harness the full potential of LLMs while ensuring patient safety and maintaining the integrity of medical practices. In this section, we attempt to identify these limitations and related safety issues in medicine.
In the field of AI, a hallucination or artificial hallucination [57] is a confident response by an AI that does not seem to be justified by its training data or real-world facts. In order to answer the request of the user, LLMs may make up responses that may not be true. Hallucination about diagnoses, treatment and patients outcomes in texts generated by LLMs will confuse and mislead decisions by both clinicians and patients. Math calculation ability of LLMs has been recently reported to be an issues in medicine. According to Barker et al [58], one in five medication doses during hospital stages are given with errors, and more than 7 million patients per year in the US are affected by such errors [59]. There, AI and LLMs hold huge opportunities to improve these issues. Recent work shows that LLMs, such as GPT-4, is able to conduct calculations or code calculator app, indicating their potential to revolutionize the way healthcare professionals approach medication dosing and reduce the likelihood of errors. However, the integration of these advanced technologies in clinical settings is not without its challenges. Prof. Kohane’s observation about GPT-4’s inaccuracies in calculating the Pearson correlation for a patient’s salt intake and systolic blood pressure serves as a critical reminder of the limitations and pitfalls associated with relying solely on AI for healthcare decisions [60]. Such discrepancies highlight the importance of incorporating an additional layer of verification, particularly for complex mathematical tasks that have direct implications for patient care.
Concerns were also raised about current LLMs’ difficulty to handle complex logic. According to Dziri et al [61], through multi-digit multiplication, logic grid puzzles, and a classic dynamic programming problem, they show empirically that LLMs cannot handle compositional tasks that require breaking down problems into compositional sub-steps and synthesizing these steps into a precise answer. In fact, transformer models solve compositional tasks by reducing multi-step compositional reasoning into linearized subgraph matching without necessarily developing systematic problem-solving skills. This will lead to unreliable results for compositional tasks in medical settings such as treatment design that requires multiple step of logic reasoning based on observation from different observations. Everyday medical operation requires strong common-sense reasoning ability of the AI models. As reported by Lee et al. [60], GPT-4 is able to have amazing performance in reasoning how to transfer different patients to different destinations and infer thoughts of the nurse who is arranging the transfer. However, some studies show limitations and instability of performance of LLMs in various common sense reasoning tasks such as in physical relation among household items[62]. If we do not have a way to estimate LLMs’ limit, and when they can be trusted, it will be very difficult to develop robust LLM based tools to handle important medical use cases such as arrangement of medical devices.
Interaction with private patient’s Information always requires the highest level of caution. As stated by ChatGPT itself, the models don’t explicitly memorize input data. However, there are still concerns about potential data leakage or the possibility of extracting specifics about the training data. In the most recently deployed LLMs, local users do not have access to the model but need to access the model by providing queries and contextual data to an online portal. In healthcare, most data belongs to individual patients and is protected by law [63].
Another challenge in LLMs is physical world understanding. When inputting the question “Can you describe how you perceive or experience the physical world?” to ChatGPT, we got the following answer: “I don’t have sensory experiences. While I can provide information about the physical world based on my training data, I don’t ‘experience’ or ‘understand’ it as humans do.” Even visual-large language models such as ChatGPT-V recently became available [64], we still lack an efficient way to enable LLMs to interact with the physical world, such as asking questions like “Which of the three ventilators in the next room should I use for the patient seen by Jone yesterday morning in his clinic?”, which limits usage of LLMs in many healthcare related activities.
Fact-checking in real-time is another problem of using LLMs in medicine. The users can not verify the real-time accuracy of statements or determine the truthfulness of recent claims, such as generated clinical summaries. The user sometimes can rely on the training data for factual information but not all the time. As many LLMs are trained offline, the lack of up-to-date information prevents the use of LLMs in healthcare operations which require frequent updates of entirely new information. For instance, the information used to train models like ChatGPT is typically a few months old. In the healthcare sector, where current information is crucial, the slow pace at which models are updated by their developers poses a significant limitation. Therefore the hospital users often needs to tune their own version of LLMs either using service from large model providers or based on open source models [65]. When interacting with LLMs, we may not always receive the outcome we want due to the limitations in communication between human users and the models. These limitations include and are not limited to LLMs acting as a echo chambers and not following instruction (see more in Table 3).
Context length is another challenge in some LLMs. Even though many LLMs and related transformer architectures have shown significant advancements in understanding and generating complex language structures, there is a notable limitation in their context length capabilities, even with recent increases in newer version such as GPT-4-omini [66, 67]. This constraint becomes especially apparent when comparing the memory of LLMs to humans, particularly across diverse conversations and documents. The limitation of accessible data types poses a another challenge for clinicians, patients, and researchers looking to leverage LLMs in medical practice or research. Currently, these models predominantly excel with textual data and images, which they can analyze and interpret with a high degree of accuracy. However, when it comes to other crucial types of data in the medical field, such as EEG waves, protein 3D structures, or genomic sequences, ordinary LLMs face limitations. These data types embody complex patterns that require not only advanced processing to convert them into a format understandable by LLMs but also necessitate the development of specialized algorithms capable of capturing the intricacies inherent to such information. Healthcare data is inherently multimodal and longitudinal, with only a fraction of data in structured forms. Rather than textual and visual data, most large generative models, including GPT-4V and Gemini, are not able to process a large number of different data modalities. Even a model is adjusted for different data modality, researchers still lack of a reliable set of benchmark to systematically evaluate model performance in medical settings as evaluating generative AI models is challenging. The vast amount of routinely collected real-world clinical data, such as electronic medical records or remote sensing data, represents a significant opportunity for the growth of multimodal AI models. Yet, the AI for medicine community still searches for effective methods to harness these multimodal real-world datasets.
Unlike many other fields, it is extremely challenging to gather a huge amount of relevant data with high quality for specific practices in healthcare, such as dentistry. Despite rigorous efforts to sanitize and filter the vast amount of training data, it takes a significant amount of efforts, if not impossible, to eliminate all harmful and inappropriate content, which may inadvertently propagate through the responses generated by LLMs. Inherently, these LLMs operate as sophisticated pattern-matching machines without a genuine understanding of the data they are trained on, which occasionally leads to nonsensical or inappropriate responses with erroneous domain knowledge [68, 69].
Table 3: Limitations of LLMs in communication.
Not Following Instruction At times, LLMs may choose not to follow instructions by the user closely. It is reported by Lee et al. that when an early version of GPT-4 was instructed to generate a document with certain patterns, it generated a different document due to a preference [60]. This will result in producing unwanted information, which leads to misleading medical decisions.
Denial on Errors In certain cases, even if the user points out the mistake by the LLMs, the AI may deny its mistake and insist on the wrong results. One such example the wrong Pearson correlation between salt intake and blood pressure until Prof. Kohane provides a math equation for Pearson correlation[60].
Sensitivity to Ambiguous Queries If a question is ambiguous or lacks context, LLMs might infer the user’s intention incorrectly and provide an answer that doesn’t align with what the user had in mind[70].
Documents in minor languages The amount of information available to train LLMs varies among languages; LLMs may perform better in one language than another which may lead to unfair performance and quality of healthcare for clinicians and patients using different languages [71].
5 Conclusion

### v2

@@ INTRODUCTION
The problem of bias in AI model performance in medicine is an important concern as it impacts the reliability and fairness of medical decisions, potentially leading to suboptimal care and health disparities among different population subgroups 59, 60, 61. The development and deployment of AI in medicine faces several bias-related challenges: (1) Label Bias: The labels in training data might reflect existing biases in clinical practice, mirroring the unequal attention and care provided to different demographic groups, thereby influencing AI outcomes. (2) Algorithmic Bias: The choice of algorithms, features, and model parameters can introduce and amplify biases. This affects model predictions and outcomes, potentially leading to inequitable healthcare delivery. (3) Evaluation Bias: Evaluation metrics and benchmarks might be biased towards certain groups or conditions. This can mask the true performance of AI models across diverse populations. (4) Socioeconomic and Cultural Bias: Differences in access to healthcare resources, healthcare-seeking behaviors, and disease prevalence across socioeconomic and cultural groups can lead to bias in construction of AI models. (5) Historical Bias: Historical inequalities and prejudices in medical research and clinical practice can be perpetuated and amplified through biased AI models, necessitating a conscientious approach to model training and validation.
While general AI models face various bias-related challenges, LLMs introduce additional bias safety concerns that warrant careful consideration in medical applications 62, 63, 64. Firstly, the expansive training datasets of LLMs may embed gender, race, societal, cultural, institutional and political biases, which can inadvertently perpetuate stereotypes or inequities in medical contexts 62, 63, 65, 66, 67, 68, 69. For instance, LLMs may demonstrate geographic bias, favoring regions with greater representation in the training data, potentially leading to inaccurate or discriminatory recommendations for underrepresented areas. Furthermore, as illustated by Schmidgall et al. 62, cognitive biases, such as self-diagnosis bias, confirmation bias, and recency bias, significantly influence the performance of LLMs in medical contexts. These biases can compromise diagnostic accuracy and decision-making, especially when models are exposed to prompts that emulate real-world scenarios involving clinical biases. The development of benchmarks like BiasMedQA emphasizes the critical need for tailored mitigation strategies for the safety of medical LLMs. In addition, LLMs exhibit implicit and source biases due to the statistical patterns inherent in their training data, which can disproportionately favor familiar or frequent inputs while underrepresenting less common but critical medical cases. Such biases can compromise the reliability of LLMs in addressing rare diseases or conditions prevalent in specific populations. Biases introduced during fine-tuning or safety interventions can also lead to unintended trade-offs, such as diminished contextual accuracy or overgeneralization in nuanced scenarios.
@@
Fine-tuning AI models to local data in medicine is essential to adapting models to the specific characteristics of new patient populations and healthcare settings. However, it poses several challenges that impact the effectiveness and generalizability of models 70, 71. Apart from the aforementioned data related challenges including Data Scarcity, Data Quality and Variability, Heterogeneity, and Data Shifting, Computational Resources may also lead to safety issues, given that the limitation of computational resources in local environments can be a significant barrier to the fine-tuning of complex models. Moreover, Domain Expertise can be a problem as well, as the absence of local expertise in fields like AI and medicine can pose challenges to the adaptation and validation of models.
In addition to these, Regulatory and Ethical Issues also pose great challenges for general AI models when applied to a new patient population, and the emergence of LLMs with their broad range of applications further amplifies the severity of this issue72, 73, 74, 75. Firstly, data privacy and security concerns are paramount, as LLMs often rely on extensive datasets that may include sensitive patient information. Without rigorous safeguards, there is a risk of unauthorized data access, re-identification of anonymized data, or breaches during model deployment to new populations. The lack of transparency around the provenance of training data further complicates compliance with intellectual property and regulatory standards. Questions about whether datasets were appropriately licensed or whether their use respects patient consent add layers of complexity. Additionally, the high variability and context-sensitivity of LLM outputs create challenges in meeting regulatory requirements for medical applications. Issues like hallucinations can have severe implications in clinical settings, such as misdiagnoses or inappropriate treatment recommendations. Moreover, existing regulations, such as the EU Artificial Intelligence Act 76, may not fully address the unique capabilities and risks of LLMs, necessitating the development of more tailored frameworks.
Alignment problems of artificial intelligence in medicine
@@
Besides performing different tasks reliably and robustly, it is important to ensure the AI models’ and LLMs’ behaviors reflect human users’ goals and values. In this section, we discuss the challenges of alignment problems for AI in medicine, with a focus on technical and normative alignment as well as AI-clinician Interaction.
@@
The alignment problem in AI refers to the challenge of ensuring that artificial intelligence systems reliably behave in ways that are in accordance with human values, intentions, and expectations. This problem arises from the inherent difficulty in specifying complex human values and objectives precisely and completely in a manner that AI systems can understand and implement. The goal of AI value alignment is to ensure that AI as well as LLMs is properly aligned with human values 77.
@@
The challenge of alignment has two parts. The first is technical, focusing on the formal incorporation of values or principles into AI to ensure their reliable and appropriate actions. Real-world instances of AI misalignment have been observed, such as chatbots promoting abusive content when interacting freely online 78. Moreover, this challenge escalates with more advanced AIs, raising issues like ‘reward-hacking’—where agents find unanticipated methods to fulfill objectives, diverging from intended outcomes—and the difficulty of evaluating AIs whose cognitive abilities might significantly surpass human capabilities 79.
@@
The second aspect of the value alignment problem is normative, raising the question of which values or principles should be encoded into AI. This can be viewed through two lenses: minimalist and maximalist conceptions of value alignment. The minimalist approach seeks to bind AI to a viable framework of human values to prevent unsafe outcomes, while the maximalist perspective aims for alignment with the most accurate or optimal human values on a broader societal or global scale. Although the minimalist perspective rightly notes that optimizing for nearly any single metric could yield adverse outcomes for humans, advancing beyond this approach might be necessary for achieving full alignment with AI. This is because AI systems could be safe and reliable but still fall short of optimal or truly desirable outcomes 77. Here we conclude the alignment challenges as in Table 2.
@@
Scalability and Complexity With increasing complexity and capability, ensuring that AI models remain aligned with human values and controlling their behavior become more challenging.
@@
While general AI models face alignment challenges, LLMs encounter additional safety concerns due to their expansive capabilities and widespread applications80, 81, 82. One significant challenge lies in fine-tuning and adapting LLMs to domain-specific safety standards, such as those outlined by the American Medical Association’s Codes and Principles of Medical Ethics 83, 84. Ensuring consistent adherence to these principles across diverse tasks and settings remains a complex and unresolved problem. Moreover, the dynamic nature of medical practice exacerbates alignment difficulties, as LLMs must remain up-to-date with evolving standards, treatments, and ethical considerations.
Another great challenge in aligning medical LLMs stems from the lack of robust tools for interpreting or explaining their behavior. Current interpretability methods often fail to provide faithful or comprehensive insights into how LLMs process inputs or generate outputs, leaving critical gaps in understanding their reasoning pathways. This opacity poses serious risks in medical applications, where harmful outputs—such as misinformation or unethical recommendations—require precise diagnosis of their root causes to enable targeted alignment interventions. For example, LLMs may generate plausible but incorrect medical advice due to unintended associations formed during training, yet without reliable interpretability tools, it becomes challenging to identify whether such behavior results from training data biases, spurious correlations, or architectural flaws.
In addition to the challenges of use and fine-tuning, evaluating the alignment of LLMs is a critical aspect of ensuring their safety and reliability, particularly in the medical domain. Evaluation methodologies face several confounding factors, such as prompt-sensitivity 85, 86, test-set contamination 87, and biases in human-based assessment frameworks 88. For example, the performance of an LLM can vary significantly based on the phrasing of the prompt, complicating the consistent measurement of alignment quality. Furthermore, targeted training and fine-tuning of LLMs can lead to overfitting on evaluation benchmarks, giving an inflated impression of alignment without addressing underlying issues.
@@
Beyond technical and normative alignment, AI-clinician interaction and alignment is also a fundamental aspect of implementing AI and LLM solutions in medical practice. A symbiotic collaboration between AI and clinicians, achieved through consensus finding or disagreements cross-validation, would significantly improve the quality of patient care and reduce clinical operation costs 89. Several challenges need to be addressed to ensure effective and seamless integration of AI and LLMs into clinical workflows. Complex and non-intuitive interfaces can impede effective interaction between clinicians and AI tools, adversely affecting adoption and user experience. Skepticism about the reliability of AI predictions may arise among clinicians, particularly when models lack transparency, interpretability, or validation in real-world settings 90. Incorporating AI tools into existing clinical workflows can be challenging, potentially causing disruptions and diminishing efficiency. Issues concerning patient privacy, data security, and accountability in cases of erroneous AI predictions can affect clinicians’ willingness to employ these tools. The performance of AI tools can be compromised by using inaccurate, incomplete, or outdated data, reducing their utility and reliability in clinical settings. A gap in AI understanding among clinicians and insufficient medical knowledge among AI developers can lead to less-than-optimal AI tool development and implementation. Inadequate training and education regarding AI tools can hinder clinicians’ ability to utilize and interpret outputs from them effectively. The integration of AI tools into clinical practice might alter the dynamics of patient-clinician relationships, potentially impacting trust, communication, and the overall quality of care. Advanced AI models, such as deep neural networks, often lack transparency, making it challenging to understand and anticipate their behaviors and underlying decision-making processes.
Building Public Confidence in the Safety of AI for Clinical Practice, Healthcare System Operation and Social Impact
In the previous sections, we highlighted that adoption of AI and LLMs in real-world medical settings remains limited due to persistent safety concerns. In this section, we explore potential strategies to foster trust and build confidence among patients, clinicians, and the public in the safety of AI within clinical practice, healthcare operations, and its broader social impact.
Confidence in AI safety for clinical practice
@@
In clinical practice, AI safety concerns exist in various forms, each carrying distinct risks. Although AI and LLMs are usually designed and functioning as assistive tools to support clinical decisions (e.g., diagnostic-AI, clinical monitoring, etc.91, 92, 93), unrobust, unsafe, and unexplainable black-box AI models could potentially result in disease misdiagnosis and subsequent mismanagement 94. It’s been known that AI models with high predictive performance can be poorly calibrated, and calibration-reporting guidelines, despite well adopted in the medical community 95, are rarely followed in the engineering research studies. Uncertainty measures of predicted outcomes, parameters, and statistics (e.g., confidence intervals) are also commonly under-reported in incomplete and/or inconsistent ways 96. Poor generalization in AI systems can lead to misdiagnoses and inappropriate treatments, impacting patient health resource allocation and eroding trust in AI technologies. Biased models exacerbate these issues, particularly affecting underrepresented groups, leading to misdiagnoses, inappropriate treatments, and widening health disparities. Lastly, the challenge of fine-tuning of AI models using local data is essential to ensure equitable healthcare outcomes, underscoring the need for models sensitive to diverse patient populations and healthcare settings. Each of these risks highlights the critical need for rigorous testing, validation, and ethical considerations in the deployment of AI in medicine in order to build confidence among the users97, 98, 99, 100.
Confidence in AI safety for healthcare system operations
@@
The integration of AI and LLMs in healthcare systems, such as hospital operations, community clinic operations, pharmacy operations and health insurance, brings various safety issues that could lead to risks and dangers. Data security and privacy is one of the largest issues when adopting AI and LLMs in healthcare operations, as AI systems require access to vast amounts of sensitive patient data. There is a risk of data breaches, unauthorized access, or misuse of data, which can lead to privacy violations and compromise patient confidentiality 101, 102. Therefore, significant improvement of data protection will be vital to build confidence and trust in using AI and LLMs in daily healthcare system operation. Another issue is robustness, which refers to the possibility that AI systems might make errors in diagnosis or treatment recommendations due to limitations in their algorithms or training data. These errors can have serious consequences for patient health 103. Ensuring that AI systems comply with existing healthcare regulations and standards is crucial. Non-compliance can lead to legal and ethical issues. Additionally, interoperability and system integration pose challenges in healthcare as integrating AI and LLMs into existing systems can be complicated due to inter-operator discrepancies and differences in data standards, which can lead to inefficiencies and errors in patient care 36. Resource allocation should also be taken into consideration, as AI-driven decisions might influence the distribution of resources in healthcare, potentially prioritizing certain groups or treatments over others, which could exacerbate inequalities 104. Apart from these technical problems, dependence and skill degradation among healthcare workers can occur due to over-reliance on AI as well as LLMs, leading to a decrease in their decision-making skills 105.
Confidence in social impacts
@@
AI safety issues in medicine when presented on a societal scale, can lead to a range of broader risks and challenges. Here we outline general issues that may arise from policy level down to population level in Table 3. These include public trust and perception, the ethical and societal norms, and the bias, fairness issues and more. To foster public confidence in the application of AI in medicine, it is essential to address these social issues with careful consideration and meticulous attention to detail.
@@
Table 3: Social risks to consider when building confidence in AI for medicine
@@
Widening Health Disparities If AI technologies are primarily accessible to well-resourced healthcare systems or populations, it could exacerbate existing health disparities. This has been observed in other technological advances and is termed the Matthew principle. Those without access to AI-enabled healthcare might receive lower quality care, widening the health outcome gap between different socio-economic groups 106.
@@
Policy and Regulatory Challenges As AI continues to advance, it may outpace current regulations and ethical guidelines, leading to legal and ethical dilemmas. Ensuring that policies keep up with technological advancements is crucial but challenging, and failure to do so could lead to societal risks 107.
@@
Misinformation and Manipulation Risks There’s a risk of AI being used to spread misinformation about health issues or being manipulated to provide misleading information, which can have widespread negative effects on public health and safety 108.
@@
Global Inequities The uneven distribution and adoption of AI in healthcare across different countries can exacerbate global health inequities. Countries with limited access to AI technology may fall behind in healthcare advancements 109.
@@
Dependency and System Vulnerabilities Over-reliance on AI systems can make healthcare infrastructure vulnerable to technical failures, cyberattacks, or other disruptions, potentially leading to widespread healthcare service interruptions 110.
CONCLUSION